One number is four more than another. Their sum is eight. Find both numbers. (Enter solutions from smallest to largest.)
Answer by ewatrrr(24785)


Hi
One number is four more than another
Let x and (x+4) represent the two numbers
Question states***
x + (x+4) = 8
Solving for x
2x = 4
x = 2
(x+4)= 6 The two numbers are 2 and 6.
CHECKING our Answer***
2 + 6 = 8